# Adaptive Interactive Navigation of Quadruped Robots using Large Language Models

### v1

@@ Abstract
Robotic navigation in complex environments remains a critical research challenge. Traditional navigation methods focus on optimal trajectory generation within free space, struggling in environments lacking viable paths to the goal, such as disaster zones or cluttered warehouses. To address this gap, we propose an adaptive interactive navigation approach that proactively interacts with environments to create feasible paths to reach originally unavailable goals. Specifically, we present a primitive tree for task planning with large language models (LLMs), facilitating effective reasoning to determine interaction objects and sequences. To ensure robust subtask execution, we adopt reinforcement learning to pre-train a comprehensive skill library containing versatile locomotion and interaction behaviors for motion planning. Furthermore, we introduce an adaptive replanning method featuring two LLM-based modules: an advisor serving as a flexible replanning trigger and an arborist for autonomous plan adjustment. Integrated with the tree structure, the replanning mechanism allows for convenient node addition and pruning, enabling rapid plan modification in unknown environments. Comprehensive simulations and experiments have demonstrated our method’s effectiveness and adaptivity in diverse scenarios. The supplementary video is available at page: https://youtu.be/W5ttPnSap2g.
Index Terms:
Interactive navigation, large language model, task and motion planning, quadruped robot.
@@ I Introduction
Navigation in diverse and complex environments is a pivotal challenge in robotics, necessitating innovative approaches to ensure effective and adaptable decision-making and planning. Recent research has demonstrated the potential of robotic platforms to traverse cluttered environments and navigate to the desired position, such as utilizing quadrupedal robots [1, 2] for maneuvering across various terrains. Despite the progresses, conventional navigation strategies primarily focus on collision-free trajectory planning within a given free workspace, which significantly limits their applicability in real-world scenarios such as disaster rescue or warehouse logistics, where collision-free paths may not exist, and the robot must interact with the environment to create viable routes.
@@
To address these limitations, previous works have explored navigation among movable obstacles (NAMO), where the robot can actively interact with and reposition movable obstacles to create viable paths. This NAMO research expands the scope of robotic navigation, enabling operation in highly cluttered and confined environments. However, some prior research [3] focused on global planning within movable obstacles but idealized interactive behavior in simplified simulations, which limited their applicability in real-world situations. To address this limitation, Simon et al. [4] have explored integrating mobile platforms with robotic arms, employing whole-body control to relocate obstacles and create pathways, and modeling more realistic interaction behaviors to enhance applicability in real-world environments. However, NAMO approaches only focus on removing obstacles without utilization of movable obstacles as tools, such as leveraging shorter obstacles as stepping stones to help the robot traverse higher hurdles impassable originally, thereby limiting their abilities to solve long-horizon interactive navigation tasks that require multi-step strategic utilization of environmental objects.
@@
Figure 1: Interactive navigation in challenging environments. In contrast to conventional obstacle-avoidance navigation systems that passively adapt to predefined free space and fail to reach goals in environments without feasible paths, our method (AINav) can push aside obstacles to create pathways within congested obstacles or utilize tools to traverse excessively high hurdles, tackling interactive navigation tasks in challenging environments.
@@
Leveraging the advanced understanding and reasoning abilities of large language models (LLMs), recent works have utilized LLMs combined with pre-trained atomic skill libraries to interact with environments to address complex long-horizon tasks [5, 6]. They formulate the complex interactive tasks as task and motion planning (TAMP) problems, where task planning involves task decomposition to assist in completing final tasks, such as selecting which objects to interact with and determining the sequence of interactions, while motion planning considers how to execute the high-level task by a controller to accomplish the task. However, these methods require global scene description from an omniscient perspective, which is often unavailable in real-world scenarios where robots operate with limited, egocentric observations. This limitation restricts their ability to adapt to dynamic and partially observable environments. Closed-loop task planning can respond to new observations through iterative replanning [7, 8]. Nevertheless, previous works commonly focus on passive failure recovery, lacking the ability to proactively utilize new information and adjust plans, which represents a critical aspect for effective navigation in partially observable environments. The primary challenge we aim to address is how to create effective plans for navigation tasks in unknown interactive environments while adaptively responding to newly acquired information.
@@
To overcome these limitations and challenges, we propose the Adaptive Interactive Navigation (AINav) approach to tackle interactive navigation tasks (Fig. 1). The proposed system demonstrates dual capabilities: it actively interacts with the environment to change the workspace for achieving navigation objectives, and adaptively reacts to new observations in unknown environments. This dual functionality enables effective navigation in unfamiliar and interactive environments using egocentric sensing information, without relying on global environmental knowledge. While recent advancements in robot parkour have demonstrated the capability to traverse across varying heights relative to the size of the quadruped robot, our research emphasizes enabling the robot to actively manipulate the environment, tackling challenges that extend beyond the scope of conventional passive obstacle avoidance and terrain adaptation navigation systems. The main contributions can be summarized as follows:
@@
  • We develop a comprehensive skill library with reinforcement learning (RL) to equip the robot with advanced locomotion and interaction capabilities. Integrated with customized reward design and task curriculum, robots can learn to navigate and manipulate robustly in complex environments.
  • We introduce an adaptive replanning mechanism including two LLM-based modules, an advisor and an arborist, where the advisor adaptively triggers replanning based on new observations, while the arborist automatically reconstructs the decision tree for optimal task execution.
We validate our method across comprehensive simulations and real-world experiments, revealing its efficacy in tackling challenging navigation tasks and its capacity for fast adaptation to new observations in unknown environments. Through this novel framework, we aim to extend the boundaries of robotic navigation, offering an effective and computationally efficient solution for interactive navigation in complex environments. The remainder of this paper is organized as follows: Section II reviews related work in robotic navigation and interactive planning. Section III presents the proposed AINav framework with specific methods. Section IV describes the experimental setup and results. Finally, Section V concludes the paper and discusses future directions.
@@ II Related Work
II-A Quadruped Robot Navigation
Quadruped robots have demonstrated agile capabilities and terrain traversability in diverse complex environments in recent researches [9, 10]. Based on the superior agility and adaptivity, quadruped robot navigation has made significant progress through various kinds of approaches, including sampling-based method [11], learning-based method [12, 13]. Recently, benefiting from the powerful understanding and reasoning ability of pre-trained large models, abundant works utilized the LLM and vision-language models (VLM) to understand the semantic environment information for more flexible terrain traverse, so as to tackle navigation tasks [14, 15]. However, current approaches often assume the existence of navigation paths, neglecting the possibility that such paths may not exist in real cluttered scenarios. Zhang et al. [16] presented an interactive navigation framework that instructs robots to navigate in an environment with traversable obstacles with LLMs. However, they aim to adapt to the environment passively rather than actively utilizing interactive objects for task completion. Wu et al. [5] and Xu et al. [6] proposed using LLMs to interact with environments and employ tools to address complex long-horizon tasks with quadruped robots more effectively. However, they assume complete scene descriptions with omniscient information and perform open-loop planning, which is impractical for navigation tasks where the robot obtains real-time egocentric observation incrementally.
@@ II-B Navigation Among Movable Obstacles
Navigation among movable obstacles refers to the problem of robot navigation that requires manipulation of the environment. Early research [3] mainly concentrated on modeling abstract and idealized interactive behaviors, overlooking the complexities inherent in real-world physical interactions. While these approaches are effective in simplified simulation settings, their failure to account for real-world physics significantly restricts their practical applicability. Recent work by Simon et al. [4] has advanced the field by integrating mobile platforms with robotic arms and utilizing whole-body control to relocate obstacles and clear pathways. This approach incorporates more realistic interaction models, improving applicability in real-world environments. Nevertheless, current NAMO approaches predominantly focus on obstacle removal and do not fully capitalize on the potential of movable obstacles, such as utilizing shorter obstacles as steps to traverse higher barriers. This limitation restricts their ability to address more complex long-horizon tasks in more intricate scenarios.
@@ II-C Closed-Loop Task Planning with LLMs
Due to superior performance in language understanding and commonsense reasoning, LLMs are widely adopted to efficiently solve closed-loop task planning, which refers to generating a series of intermediate steps to achieve the specific goal while modifying the plan based on new observations. Many works incorporate an iterative manner for closed-loop task planning, which replans based on the description or analysis of real-time observations [17, 18]. However, these methods typically generate a single plan for execution, which often contains errors or infeasible decomposition due to model uncertainties. This increases the need for replanning and hinders computationally efficient planning. Tree-based structure for LLM reasoning and planning enhances the solution quality by generating multiple plans and select the optimal tree path with state evaluation [19, 7, 20]. However, these methods either lack a replanning mechanism or trigger replanning solely upon encountering failures and concentrate primarily on error correction, overlooking a comprehensive understanding and effective utilization of interactive environments.
III Adaptive Interactive Navigation
@@
Our framework consists of three primary components: LLM-based task planning, motion planning with pre-trained skills, and adaptive replanning, as shown in Fig. 2. The robot first captures the first-person perspective images and extracts scene understanding and object descriptions using a VLM and object detector, respectively. By combining user instructions with environmental information as inputs, the task planning module uses the LLM to construct a primitive tree for reasonable task decomposition and convenient adjustments based on new information, enabling the robot to obtain the optimal high-level primitive skeleton. Leveraging a pre-trained skill library consisting of robust locomotion and interaction skills with RL, the motion planning module executes specific low-level actions following the skeleton to complete the task. Finally, a novel adaptive replanning mechanism interprets new observations from an egocentric perspective to determine whether replanning is necessary, allowing for swift reaction to incremental environmental observations.
III-A Task Planning with Primitive Skill Tree
@@
We propose the primitive skill tree for task planning. The primitive skill tree encompasses two LLM-based roles: a proposer for primitive tree construction, and an evaluator for node evaluation and skeleton selection.
@@
Proposer: object-level to skill-level. Given the user instructions $I$ and environmental information $O$, we first utilize the LLM to propose multiple potential object-level planning strategies $P_{obj}$, where each plan specifies which objects to use and how to leverage them to achieve the task
$\begin{split}P_{obj}=\text{LLM}(I,O).\end{split}$ (1)
@@
Building upon object-level plans, we subsequently generate standardized skill-level plans $P_{skill}$, where each step corresponds to a primitive skill from the robot’s skill library $S$.
$\begin{split}P_{skill}=\text{LLM}(P_{obj},S).\end{split}$ (2)
@@
After skill-level plans are generated, we construct the primitive skill tree by merging steps of plans in $P_{skill}$ with identical historical traces. This approach offers two key advantages: First, unlike other approaches that use the LLM to directly generate a single task plan, multiple alternative plans proposed by the LLM are more robust to unexpected errors, such as misunderstandings by the LLM, limited scene observation, and environmental uncertainties. Second, compared to directly using the LLM to generate detailed plans that may lead to illogical or inexecutable plans, transitioning from holistic object utilization to specific skill planning tends to propose more reasonable and feasible object interaction strategies.
@@
Evaluator: short-term to long-term. To evaluate the proposed plans, we utilize another LLM as an evaluator to assess the immediate reward $r(n)$ of each node $n$ in the primitive tree, which considers the difficulty of primitive skills and the contribution to the specific task. For example, if the robot interacts with objects like pushing movable obstacles to create a free path for navigation, this skill is essential to completing the task, but the interaction is also difficult to complete. If one skill is non-executable, the subsequent nodes including this node will be deleted to ensure the feasibility of the plan. If a trajectory in the primitive tree reaches the target point at the end, an additional bonus is given to the leaf node as the terminal reward. Finally, we apply a backup method to update the long-term cumulative reward $Q(n)$ of nodes with
@@
where $\gamma$ is the discounted factor. By iteratively selecting the child nodes with maximal cumulative reward, we derive the optimal high-level primitive skeleton. This approach also allows for the existence of alternative strategies, enabling rapid replanning upon receiving new environmental information, which will be specifically addressed in Section III-C.
III-B Skill Pre-training with Reinforcement Learning
To robustly execute the decomposed tasks, we train a primitive skill library using RL for motion planning. The skill library includes both low-level locomotion skills, such as walking and climbing, and high-level skills like navigation and object-pushing strategies. We first train locomotion policies to develop fundamental movement capabilities on diverse terrains. Building upon the locomotion policies, we employ a hierarchical RL framework to train more advanced strategies, such as navigation policy for collision-free movement in cluttered environments and pushing policy for moving objects to designated positions.

### v2

# AINav: Interactive Navigation via LLM-based Primitive Skill Tree and Adaptive Replanning

@@ Abstract
Robotic navigation in complex environments remains a critical research challenge. Traditional navigation methods focus on optimal trajectory generation within fixed free workspace, therefore struggling in environments lacking viable paths to the goal, such as disaster zones or cluttered warehouses. To address this problem, we propose AINav, an adaptive interactive navigation approach that proactively interacts with environments to create feasible paths to achieve originally unreachable goals. Specifically, we present a primitive skill tree for task planning with large language models (LLMs), facilitating effective reasoning to determine interaction objects and sequences. To ensure robust subtask execution, we adopt reinforcement learning to pre-train a comprehensive skill library containing versatile locomotion and interaction behaviors for motion planning. Furthermore, we introduce an adaptive replanning approach featuring two LLM-based modules: an advisor serving as a flexible replanning trigger and an arborist for autonomous plan adjustment. Integrated with the tree structure, the replanning mechanism allows for convenient node addition and pruning, enabling rapid plan adaptation in a priori unknown environments. Comprehensive simulations and experiments have demonstrated AINav’s effectiveness and adaptivity in diverse scenarios. The supplementary video is available at: https://youtu.be/CjXm5KFx9AI.
@@ I Introduction
Navigation in diverse and complex environments is a pivotal problem in robotics, necessitating innovative approaches to ensure effective and adaptable decision-making and planning. Recent research has demonstrated the potential of quadruped robots to traverse cluttered environments and navigate to the desired position on complex terrains [hoeller2024anymal, han2024lifelike]. Despite the progresses, conventional navigation strategies primarily focus on optimal trajectory planning within a given free workspace, which significantly limits their applicability in real-world scenarios such as disaster rescue or warehouse logistics, where feasible paths may not exist a priori, and the robot needs to interact with the environment to create viable routes.
@@
To tackle this challenge, previous works have explored strategies for a special case, namely navigation among movable obstacles (NAMO), where the robot can actively interact with movable obstacles to create viable paths. However, prior research [muguira2023visibility] primarily focused on global planning and assumed deterministic object dynamics during interaction in simplified simulations, which limited their applicability in real-world situations. To address this limitation, Simon et al. [armleder2024tactile] have explored integrating mobile platforms with robotic arms, employing whole-body control to relocate obstacles and create pathways, and modeling more realistic interaction behaviors to enhance applicability in real-world environments. However, NAMO approaches only focus on removing obstacles without utilization of movable obstacles as tools, such as leveraging shorter obstacles as stepping stones to help the robot traverse originally impassable hurdles, thereby limiting their abilities to solve long-horizon interactive navigation tasks that require multi-step strategic utilization of environmental objects.
@@
Figure 1: Interactive navigation in challenging environments. In contrast to conventional obstacle-avoidance navigation systems that passively adapt to predefined free space and fail to reach goals in environments without feasible paths, AINav can push aside obstacles to create pathways within congested obstacles or utilize obstacles as tools to traverse excessively high hurdles, enabling interactive navigation in challenging environments.
@@
Leveraging the advanced understanding and reasoning abilities of large language models (LLMs), recent works have combined LLMs with pre-trained atomic skill libraries to interact with environments to address complex long-horizon tasks [ouyang2024long, xu2023creative]. Specifically, the interactive tasks are formulated as task and motion planning (TAMP) problems, where task planning involves task decomposition to assist in completing final tasks, such as selecting which objects to interact with and determining the sequence of interactions, while motion planning designs controllers to execute the high-level tasks. Existing TAMP methods usually require global scene description from an omniscient perspective, which is often unavailable in real-world scenarios where robots operate with limited, egocentric observations. This limitation restricts their ability to adapt to dynamic and partially observable environments. Closed-loop task planning can respond to new observations through iterative replanning [hu2023tree, wang2023describe]. Nevertheless, previous works commonly focus on passive failure recovery, lacking the ability to proactively utilize new information and adjust plans, which represents a critical aspect for effective navigation in partially observable environments. Therefore, the primary challenge we aim to address is how to create effective plans for navigation tasks in unknown interactive environments while adaptively responding to newly acquired information.
@@
To overcome these challenges, we propose the Adaptive Interactive Navigation (AINav), a closed-loop system that demonstrates dual capabilities: it actively interacts with the environment to change the workspace for achieving navigation objectives, and adaptively reacts to new observations (Fig.˜1). This dual functionality enables effective navigation in unfamiliar and interactive environments using egocentric sensing information, without relying on global environmental knowledge. While recent advancements in robot parkour [cheng2024extreme] have demonstrated the capability to traverse across varying heights relative to the size of the quadruped robot, our research emphasizes enabling the robot to actively manipulate the environment, tackling challenges that extend beyond the scope of conventional passive obstacle avoidance and terrain adaptation navigation systems. The main contributions can be summarized as follows:
@@
  • We develop a comprehensive skill library with reinforcement learning (RL) to equip the robot with advanced locomotion and interaction capabilities. Integrated with customized reward design and task curriculum, robots can learn to navigate and interact robustly in complex environments.
  • We introduce an adaptive replanning mechanism including two LLM-based agents, an advisor and an arborist, where the advisor adaptively triggers replanning based on new observations, and the arborist automatically adjusts the decision tree for optimal task execution.
We validate AINav across various simulations and real-world experiments, revealing its efficacy in tackling challenging navigation tasks and its capacity for fast adaptation to new observations in unknown environments. Through this novel framework, we aim to extend the boundaries of robotic navigation, offering an effective and computationally efficient solution for interactive navigation in complex environments. The remainder of this paper is organized as follows: Section II reviews related work in robotic navigation and interactive planning. Section III introduces the problem formulation of interactive navigation. Section IV presents the proposed AINav framework. Section V describes the experimental setup and results. Finally, Section VI concludes the manuscript and discusses future directions.
@@ II Related Work
II-A Robot Navigation in Complex Environments
Navigation in complex environments requires robots to traverse cluttered spaces, dynamically avoid obstacles, and adapt to irregular terrains [hoeller2024anymal]. Leveraging the superior agility and adaptivity of quadruped robots, significant progress has been made to complete navigation tasks across diverse complex terrain, including sampling-based methods [wellhausen2023artplanner], learning-based methods [rudin2022advanced]. Recently, benefiting from the powerful understanding and reasoning ability of pre-trained large models, abundant works utilized the LLM and vision-language models (VLM) to understand the semantic environment information for more flexible terrain traverse, so as to tackle navigation tasks [zhu2024cross, shah2023vint]. However, current approaches often assume the existence of navigation paths, neglecting the possibility that such paths may not exist in real cluttered scenarios. Zhang et al. [zhang2024interactive] presented an interactive navigation framework that instructs robots to navigate in an environment with traversable obstacles with LLMs. However, they aim to adapt to the environment passively rather than actively utilizing interactive objects for task completion. Wu et al. [ouyang2024long] and Xu et al. [xu2023creative] proposed using LLMs to interact with environments and employ tools to address complex long-horizon tasks with quadruped robots more effectively. However, they assume complete scene descriptions with omniscient information and perform open-loop planning, which is impractical for navigation tasks where the robot obtains egocentric observation incrementally.
@@ II-B Navigation Among Movable Obstacles
Navigation among movable obstacles refers to the problem of robot navigation that requires manipulation of the environment. Early research [muguira2023visibility] mainly concentrated on modeling abstract and idealized interactive behaviors, overlooking the complexities inherent in real-world physical interactions. While these approaches are effective in simplified simulation settings, their failure to account for real-world physics significantly restricts their practical applicability. Recent work by Simon et al. [armleder2024tactile] has advanced the field by integrating mobile platforms with robotic arms and utilizing whole-body control to relocate obstacles and clear pathways. This approach incorporates more realistic interaction models, improving applicability in real-world environments. Nevertheless, current NAMO approaches predominantly focus on obstacle removal and do not fully capitalize on the potential of movable obstacles, such as utilizing shorter obstacles as steps to traverse higher barriers. This limitation restricts their ability to address more complex long-horizon tasks in more intricate scenarios.
@@ II-C Closed-Loop Task Planning with LLMs
Due to superior performance in language understanding and commonsense reasoning, LLMs are widely adopted to efficiently solve closed-loop task planning, which refers to generating a series of intermediate steps to achieve the specific goal while modifying the plan based on new observations. Many works incorporate an iterative manner for closed-loop task planning, which replans based on the description or analysis of real-time observations [liu2023reflect, wang2024llm]. However, these methods typically generate a single plan for execution, which often contains errors or infeasible decomposition due to model uncertainties. This increases the need for replanning and hinders computationally efficient planning. Tree-based structure for LLM reasoning and planning enhances the solution quality by generating multiple plans and select the optimal tree path with state evaluation [yao2024tree, hu2023tree, zhao2024large]. LLM-BT [zhou2024llm] combines LLM with behavior trees to automatically update the tree, addressing the reactive rigidity of traditional methods to adapt to environmental changes. However, these methods either lack a replanning mechanism or trigger replanning solely upon encountering failures and concentrate primarily on error correction, overlooking a comprehensive understanding and effective utilization of interactive environments.
Figure 3: Task planning with the proposer and the evaluator.
III Problem Formulation
We formalize interactive navigation as a 6-tuple TAMP problem $(\mathcal{X}_{init},\mathcal{X}_{goal},\mathcal{O}_{s},\mathcal{O}_{m},\mathcal{A},\mathcal{C})$, with robot’s initial position $\mathcal{X}_{init}$, navigation goal position $\mathcal{X}_{goal}$, fixed obstacles $\mathcal{O}_{f}$, movable obstacles $\mathcal{O}_{m}$, action space $\mathcal{A}$, and constraints $\mathcal{C}$. The objective of interactive navigation is to execute a sequence of actions $\pi=[a_{1},\ldots,a_{n}]$ that enables the robot to move from initial position $\mathcal{X}_{init}$ to goal position $\mathcal{X}_{goal}$, where each action $a_{i}=(p_{i},q_{i})\in\mathcal{A}$ ($i=1,\ldots,n$) consists of a primitive skill $p_{i}$ from the pre-defined skill library and continuous parameters $q_{i}$ for execution policy. Finally, each action $a_{i}$ is executed by a pre-trained policy that generates the joint control commands for quadruped robot actuation. We assume that the robot knows whether the objects within its field of view are movable, i.e., whether an object belongs to $\mathcal{O}_{f}$ or $\mathcal{O}_{m}$. Unlike traditional navigation problems that assume a fixed free space and typically the existence of a feasible path to the goal, interactive navigation tasks consider scenarios where no feasible path exists in the original workspace, and the robot needs to proactively change the workspace through interactions to create feasible paths. We propose a hierarchical planning framework, AINav, which decides the primitive skills to execute for task planning and specifies execution details for motion planning, jointly optimizing discrete skill selection and continuous execution parameters to tackle interactive navigation problems.
IV Adaptive Interactive Navigation
IV-A Framework Overview
@@
AINav consists of three primary components: LLM-based task planning, motion planning with pre-trained skill library, and adaptive replanning, as shown in Fig.˜2. The robot first captures the first-person perspective images, and extracts scene understanding (e.g., whether pathways are obstructed) using a VLM, and an object detector provides object descriptions including spatial positions and geometric properties. By combining user instructions, including navigation goal, skill library, and constraints, with environmental information as inputs, the task planning module uses the LLM to construct a primitive skill tree for reasoning and task decomposition, enabling the robot to obtain the optimal high-level primitive skeleton. Taking the height data from lidar as input, the motion planning module executes specific low-level actions following the skeleton by leveraging a pre-trained skill library consisting of robust locomotion and interaction skills. Finally, a novel adaptive replanning mechanism interprets new egocentric observations, not only determining whether replanning is necessary but also adjusting the current plan for swift adaptation to incremental environmental observations. The detailed prompt of the proposed LLM-based proposer, evaluator, advisor, and arborist can be found in Section˜-A.
IV-B Task Planning with Primitive Skill Tree
@@
We propose the primitive skill tree for task planning shown in Fig.˜3. The primitive skill tree encompasses two LLM-based roles: a proposer for primitive tree construction denoted as $\text{LLM}_{p}$, and an evaluator for node evaluation and skeleton selection, denoted as $\text{LLM}_{e}$.
@@
Proposer: object-level to skill-level. Given the user instructions $I$ and environmental information $O$, we first utilize $\text{LLM}_{p}$ to propose multiple potential object-level strategies $P_{obj}$, where each plan specifies which objects to interact with and how to leverage them to achieve the navigation goal,
$\begin{split}P_{obj}=\text{LLM}_{p}(I,O).\end{split}$ (1)
@@
Building upon object-level plans, we subsequently generate standardized skill-level plans $P_{skill}$, where each step corresponds to a primitive skill from the robot’s skill library $S$,
$\begin{split}P_{skill}=\text{LLM}_{p}(P_{obj},S).\end{split}$ (2)
@@
Each primitive skill consists of the skill name and symbolic parameters, including walk (target position), climb (target position), navigate (target position), and push (object id, target position). After skill-level plans are generated, we construct the primitive skill tree by merging steps of plans in $P_{skill}$ with identical historical traces. Specifically, each plan path is iteratively added to the tree. If a sequence of primitive skills and parameters already exists in the tree, new plans are merged with the existing structure at that common node. Otherwise, a new tree branch is created. This approach offers two key advantages: First, unlike other approaches that use the LLM to directly generate a single task plan, multiple alternative plans proposed by the LLM are more robust to unexpected situations, such as misunderstandings by the LLM, limited scene observation, and environmental uncertainties. Second, compared to directly using the LLM to generate detailed plans that may lead to illogical or inexecutable plans, transitioning from holistic object reasoning and utilization to specific skill planning tends to propose more reasonable and feasible interaction strategies.
@@
Evaluator: short-term to long-term. To evaluate the proposed plans, we utilize $\text{LLM}_{e}$ to assess the immediate reward $r(n)$ of each node $n$ in the primitive tree generated by proposer $\text{LLM}_{p}$, which considers the contribution to the specific task and the difficulty of pre-defined primitive skills. For example, if the robot interacts with objects like pushing movable obstacles to create a free path for navigation, this skill is essential to completing the task, but the interaction is also difficult to complete. If a skill is non-executable, the subsequent nodes including this node will be deleted to ensure the feasibility of the plan. If a trajectory in the primitive tree reaches the navigation goal at the end, an additional bonus is given to the leaf node as the terminal reward. Finally, we apply a backpropagation method to update the long-term cumulative reward $Q(n)$ of nodes as follows,
@@
where $\gamma$ is the discount factor. By iteratively selecting the child nodes with maximal cumulative reward, we derive the optimal high-level primitive skeleton. The symbolic parameters are then converted into precise coordinates using the LLM for execution, which is detailed in Section˜-A. This tree-structured approach also allows the existence of alternative strategies, enabling rapid replanning upon receiving new environmental information, which will be described in Section˜IV-D.
IV-C Skill Pre-training with Reinforcement Learning
To robustly execute the decomposed tasks, we train a primitive skill library using RL for motion planning. The skill library includes both low-level locomotion skills, such as walking and climbing, and high-level skills like navigation and object-pushing strategies.